# Modeling Multistability and Hysteresis in Urban Traffic Dynamics

### v1

@@ Abstract
Discussion— In summary, we empirically analyzed the propagation and dissipation of local congestion using real-world traffic data, and observed that the recovery mechanism of individual roads is not spontaneous but depending on the states of their neighboring roads, which previous studies have usually either overlooked or simplified as spontaneous. Based on this observation in the dynamics of local congestion, we suggested a simple model of congestion propagation and dissipation, which is the modified SIS model with nonlinearly decreasing recovery rate as a function of flow states of neighboring roads. Our simple model provided qualitative prediction and explanation of the macroscopic functional phases, multistability, and hysteresis of urban road networks as a result of the interplay between the propagation and recovery rates. Such hindered recovery is not limited to traffic congestion but is also related to other collective failures in networks, especially those involving load sharing for local recovery (e.g., power system [31, 32, 33, 34], internet traffic [35, 36] and logistics network [37, 38]), which is commonly observed in social infrastructure. Therefore, we hope that our theoretical framework with a complex recovery process can provide a new tool to better understand the emergent vulnerability and resilience of these networked systems.
@@
Complex contagion has been highlighted in recent years for its ability to explain social phenomena [39, 40, 41, 42, 43]. In particular, existing studies have focused on distinguishing between the propagation processes of simple and complex contagion [44, 45, 21, 46]. We have shown that complex recovery processes in urban congestion dynamics play an important role in the hysteresis and multistability of the system, similar to the role of complex contagion in other studies. A comprehensive understanding of network dynamics can be obtained when considering not only activation but also deactivation process.

### v2

@@ Abstract
Discussion— In summary, we empirically analyzed the propagation and dissipation of local congestion using real-world traffic data, and observed that the recovery mechanism of individual roads is not spontaneous but depending on the states of their neighboring roads, which previous studies have usually either overlooked or simplified as spontaneous. Based on this observation in the dynamics of local congestion, we suggested a simple model of congestion propagation and dissipation, which is the modified SIS model with nonlinearly decreasing recovery rate as a function of flow states of neighboring roads. Our simple model provided qualitative prediction and explanation of the macroscopic functional phases, multistability, and hysteresis of urban road networks as a result of the interplay between the propagation and recovery rates. Such hindered recovery is not limited to traffic congestion but is also related to other collective failures in networks, especially those involving load sharing for local recovery (e.g., power system [albert2004StructuralVulnerabilityNorth, dobson2007ComplexSystemsAnalysis, wang2009ModelCascadingFailures, nesti2020EmergenceScaleFreeBlackout], internet traffic [fu2021ModelingAnalyzingCascading, yin2023AnalysisCascadingFailures] and logistics network [yang2020ControllabilityRobustnessCascading, chen2022ResilienceLogisticsNetwork]), which is commonly observed in social infrastructure. Therefore, we hope that our theoretical framework with a complex recovery process can provide a new tool to better understand the emergent vulnerability and resilience of these networked systems.
@@
Complex contagion has been highlighted in recent years for its ability to explain social phenomena [centola2007ComplexContagionsWeakness, karsai2014ComplexContagionProcess, monsted2017EvidenceComplexContagion, sprague2017EvidenceComplexContagion, iacopini2019SimplicialModelsSocial]. In particular, existing studies have focused on distinguishing between the propagation processes of simple and complex contagion [czaplicka2016CompetitionSimpleComplex, min2018CompetingContagionProcesses, cencetti2023DistinguishingSimpleComplex, horsevad2022TransitionSimpleComplex]. We have shown that complex recovery processes in urban congestion dynamics play an important role in the hysteresis and multistability of the system, similar to the role of complex contagion in other studies. A comprehensive understanding of network dynamics can be obtained when considering not only activation but also deactivation process.